OpenStudy (anonymous):

What is the x-intercept of 3x + 2y = 15? Select one: a. (7.5, 0) b. (0, 7.5) c. (5, 0) d. (0, 5)

OpenStudy (anonymous):

@ChrisS

OpenStudy (chriss):

At the x intercept, y will equal 0. So substitute 0 for y in the given equation, then solve for x.

OpenStudy (chriss):

3x+2y=15 3x+2(0)=15 3x=15

OpenStudy (anonymous):

its 3

OpenStudy (chriss):

3x=15 divide both sides by three \[\frac{ 3x }{ 3 }=\frac{ 15 }{ 3 }\] x=5

OpenStudy (anonymous):

so the answer is c

OpenStudy (chriss):

yes

OpenStudy (anonymous):

thnx